Diffusion Morning Report 05/29/12

Ted Berg tberg at stanford.edu
Tue May 29 06:55:10 PDT 2012


Good Morning All,
     Looking good this morning. POCl3 is down again for a leak at the 
rear fitting.
     Leftovers: Cooing water on Epi-fittings and hose are here working 
on replacement, Tylan burnbox -altering recipes TEOS and LTO are back 
up,Thermco 3 and 4 no TLC
     Resolved: Silane for Epi and ICP is back on, TEOS OK to run> No 
other major issues at this time.

					
	Date: 	5/29/12 	THICKNESS-Change at 		Priority
Tube Name/Number 	Status 				
Tylan 1 	*up* 	** 	** 	** 	**
Tylan 2 	*up* 	** 	** 	** 	**
Tylan 3 	*up* 	** 	** 	** 	**
Tylan 4 	*up* 	** 	** 	** 	**
Tylan 5-BBr3 	*up* 	** 	** 	** 	**
Tylan 6-POCl3 	*down* 	*leak again* 	** 	** 	*1*
Tylan 7-TEOS2 	*up* 	** 	*14.39/30* 	** 	**
Tylan 8-SiGe 	*down* 	*burn box* 	*3.4/30* 	** 	**
Tylan 9-Poly 	*down* 	*burn box* 	*23.7/30* 	** 	**
Tylan 10-Nitride 	*down* 	*burn box* 	*17.1/30* 	** 	**
Tylan 11-LTO 	*yellow* 	** 	*9.7./30* 	** 	**
Tylan 12-FGA 	*up* 	** 	** 	** 	**
Thermco 1 	*up* 	** 	** 	** 	**
Thermco 2 	*up* 	** 	** 	** 	**
Thermco 3 	*yellow* 	*testing TLC* 	** 	** 	**
Thermco4 	*yellow* 	*bad source controller-no TLC* 	** 	** 	**
Thermco Poly1 	*up* 	** 	** 	*
* 	**
Thermco Poly2 	*up* 	** 	** 	** 	**
Thermco Nitride1 	*up* 	** 	** 	** 	**
Thermco bpsg 	*up* 	** 	** 	** 	**
AMAT Epi Cham. A 	*down* 	*SiH4 issues-heater contactor issues* 	** 
** 	*2*
AMAT Epi Cham. B 	*down* 	** 	** 	** 	**
TEL Spa 	*yellow* 	*needs test* 	** 	** 	**
Burn Box 	*yellow* 	*testing and modifying cvd recipes* 	** 	** 	*3*
SEM 4160 	*up* 	** 	** 	** 	**
FGA2 	*up* 	** 	** 	** 	**
SiH4 for furnaces 	*on* 	*crossover ok* 	*changeover working well* 	** 	**


-------------- next part --------------
An HTML attachment was scrubbed...
URL: <http://snf.stanford.edu/pipermail/diffusion/attachments/20120529/29b12d2a/attachment.html>


More information about the diffusion mailing list